OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 80 on her algebra tests. What is the minimum score she must receive on the fifth test to have an overall test score average of at least 82? (Hint: The average of a list of numbers is their sum divided by the number of numbers in the list.)

OpenStudy (anonymous):

would it be 70

OpenStudy (anonymous):

(85 + 88 + 87 + 80 + x) / 5 = 82 (340 + x)/5 = 82 (multiply equation by 5) 340 + x = 410 x = 410 - 340 x = 70 you got it....have to make at least 70